Use the following functions that show quantity demanded and quantity supplied to answer the question. Qd = 100 – 5P and Qs = –20 + 3P. If the price is R10, then there will be a shortage of   

Mark for Review What's This


1
Expert's answer
2021-06-03T09:11:55-0400
"Qd=100-5\times10=50"

"Qs=-20+3\times10=10"

"50-10=40"